# VinciCoder: Unifying Multimodal Code Generation via Coarse-to-fine Visual Reinforcement Learning

### v1

@@ 4.1 Implementation Details
We conduct SFT on this dataset for one epoch, using Qwen2.5-VL-7B-Instruct [2] and Qwen3-VL-8B-Instruct as base models. The training is performed on 24 H800 GPUs with a global batch size of 96. In the RL stage, we utilize GRPO to train the SFT-model , allocating 16 GPUs for the policy model and 4 GPUs for reward scoring, with a global batch size of 256. The DINOv2-L is utilized for visual embedding. The SFT and RL training phases are implemented using the ms-swift [55]and Easyr1 [33] frameworks.
@@ 4.2 Evaluation Settings
We conduct comprehensive experiments to evaluate the unified multimodal code generation capabilities of VinciCoder and baseline models, including closed-source [10, 1, 25] and open-source models [2, 57, 40]. Our evaluation spans five domains, four of which utilize established benchmarks: ChartMimic [45] direct generation for chart-to-code, Design2Code [34] for webpage-to-HTML, UniSVG [21] ISVGEN set for image-to-SVG, and Image2Struct [27] for image-to-LaTeX generation. For the UniSVG, the final score is a weighted combination of two components. The low-level score, which is the average of SSIM and (1 - LPIPS), contributes 40% to the final score, while the high-level score contributes the remaining 60%. For the Image2Struct, we utilize Earth Mover Similarity (EMS) in their GitHub repository for evaluation. For the molecule-to-code task, we evaluate performance on the Cosyn-400k [49] chemistry test set, which consists of 128 molecular images paired with their ground-truth SMILES strings. We report two metrics: the execution rate and the average Tanimoto similarity (Tani. Sim.) of the generated SMILES.
@@ 4.3 Main Results
The comparison between SFT and the final models further validate the effectiveness of our two-stage training process. First, the SFT stage alone yields substantial gains, with VinciCoder-SFT significantly outperforming its base model across nearly all benchmarks. This result underscores the high quality and efficacy of our large-scale SFT dataset. The subsequent ViRL stage further enhances performance across almost all metrics, particularly visual similarity and execution rate. This improvement stems from two distinct mechanisms. The execution rate increases because the RL penalizes invalid outputs by assigning a zero reward to non-executable code. Concurrently, the ViRL strategy drives the enhanced visual similarity by directly optimizing the perceptual alignment between rendered and target images. We visualize the reward progressions during the ViRL training stage in Figure 4.
In addition to generalist models, we conduct a comparative analysis against task-specific baselines in the Appendix.
@@ 4.4 Ablation Studies
Reward Functions. We conduct several ablation studies to justify the design of our visual reward, which combines a coarse-to-fine strategy with DINOv2-L embeddings. All studies use VinciCoder-SFT-7B as the base model. First, we investigate alternative reward formulations, comparing our approach against a low-level perceptual metric (SSIM) and a semantic-level reward from the image-text aligned ViT (InternViT-300M-448px-V2.5 [9]). We select InternViT as its parameter count is the same as DINOv2-L, which allows for a controlled comparison. In addition, we further ablate the individual components of our coarse-to-fine reward strategy to isolate their impact. The results in Table 4 confirm that our proposed coarse-to-fine reward achieves the most balanced and effective performance. Our analysis indicates that self-supervised DINOv2 excel at capturing the fine-grained features essential for visual similarity evaluation, in contrast to image-text aligned ViTs, which are optimized for global semantic alignment. The ablation study also highlights the critical role of our coarse-to-fine strategy for high-resolution images. The performance degrades significantly without both reward components on the ChartMimic and Design2Code benchmarks. Conversely, performance on the lower-resolution UniSVG benchmark is stable, as its images are small enough that our method naturally reduces to a single global comparison, making the coarse and fine-grained rewards equivalent.
@@ 5 Conclusion
We introduce VinciCoder, a unified multimodal code generation model trained via a two-stage SFT-ViRL strategy in this work. We construct a large-scale dataset of 1.6M samples from various domains to perform SFT. For the subsequent RL stage, we propose a novel coarse-to-fine ViRL training strategy, which calculates visual similarity rewards from both coarse-grained and fine-grained image patches. Experimental results demonstrate the superior performance of VinciCoder, validating the effectiveness of our proposed unified multimodal code generation model and the ViRL training paradigm.
@@ 5.2 ChemDraw Metrics
6 More Experiment Results
@@ 6.1 Comparison with task-specific models
We compare VinciCoder with other domain-specific models, like chart-to-code and image-to-SVG models. The results in Table 6 show that VinciCoder performs comparably to specialized models such as MSRL [5] and ChartMaster [37] on the ChartMimic dataset. More domain-specific models are compared in the Appendix. Moreover, its performance on UniSVG exceeds that of fine-tuned specialist models, which validates the effectiveness of our unified framework for multimodal code generation.
@@
ChartCoder [54] 91.4 72.5 74.0
Chart2Code[52] 62.1 42.9 33.3
MSRL [5] 96.5 78.6 83.8
ChartMaster [37] 93.8 78.2 85.1
@@
Table 7: Comparing with task-specific image-to-SVG models. The metrics of finetuned models are from UniSVG.
@@ 7 Evaluation Prompts
ChemDraw Benchmark Prompt Instruction: Create a python script that displays a visual representation of the molecule shown in the image. Utilize ’smiles = <<SMILES>> string’ to define the smile and RKiT or Indigo for plot. Image:Refer to caption

### v2

@@ 4.1 Implementation Details
We conduct SFT on this dataset for one epoch, using Qwen2.5-VL-7B-Instruct [bai2025qwen2] and Qwen3-VL-8B-Instruct as base models. The training is performed on 24 H800 GPUs with a global batch size of 96. In the RL stage, we utilize GRPO to train the SFT-model, allocating 16 GPUs for the policy model and 4 GPUs for reward scoring, with a global batch size of 256. The DINOv2-L is utilized for visual embedding. The SFT and RL training phases are implemented using the ms-swift [zhao2025swift] and Easyr1 [sheng2024hybridflow].
@@ 4.2 Evaluation Settings
We conduct comprehensive experiments to evaluate the unified multimodal code generation capabilities of VinciCoder and baseline models, including closed-source [comanici2025gemini, anthropic_claude_4_5_sonnet, openai_gpt_5] and open-source models [bai2025qwen2, zhu2025internvl3, wang2025internvl3]. Our evaluation spans five domains, four of which utilize established benchmarks: ChartMimic [yang2024chartmimic] direct generation for chart-to-code, Design2Code [si2024design2code] for webpage-to-HTML, UniSVG [li2025unisvg] ISVGEN set for image-to-SVG, and Image2Struct [roberts2024image2struct] for image-to-LaTeX generation. For the UniSVG, the final score is a weighted combination of two components. The low-level score, which is the average of SSIM and (1 - LPIPS), contributes 40% to the final score, while the high-level score contributes the remaining 60%. For the Image2Struct, we utilize Earth Mover Similarity (EMS) in their GitHub repository for evaluation. For the molecule-to-code task, we evaluate performance on the Cosyn-400k [yang2025scaling] chemistry test set, which consists of 128 molecular images paired with their ground-truth SMILES strings. We report two metrics: the execution rate and the average Tanimoto similarity (Tani. Sim.) of the generated SMILES.
@@ 4.3 Main Results
The comparison between SFT and the final models further validate the effectiveness of our two-stage training process. First, the SFT stage alone yields substantial gains, with VinciCoder-SFT significantly outperforming its base model across nearly all benchmarks. This result underscores the high quality and efficacy of our large-scale SFT dataset. The subsequent ViRL stage further enhances performance across almost all metrics, particularly visual similarity and execution rate. This improvement stems from two distinct mechanisms. The execution rate increases because the RL penalizes invalid outputs by assigning a zero reward to non-executable code. Concurrently, the ViRL strategy drives the enhanced visual similarity by directly optimizing the perceptual alignment between rendered and target images. We visualize the reward progressions during the ViRL training stage in Figure 4. In addition to generalist models, we conduct a comparative analysis against task-specific baselines in the Appendix.
@@ 4.4 Ablation Studies
Reward Functions. We conduct several ablation studies to justify the design of our visual reward, which combines a coarse-to-fine strategy with DINOv2-L embeddings. All studies use VinciCoder-SFT-7B as the base model. First, we investigate alternative reward formulations, comparing our approach against a low-level perceptual metric (SSIM) and a semantic-level reward from the image-text aligned ViT (InternViT-300M-448px-V2.5 [chen2024expanding]). We select InternViT as its parameter count is the same as DINOv2-L, which allows for a controlled comparison. In addition, we further ablate the individual components of our coarse-to-fine reward strategy to isolate their impact. The results in Table 4 confirm that our proposed coarse-to-fine reward achieves the most balanced and effective performance. Our analysis indicates that self-supervised DINOv2 excel at capturing the fine-grained features essential for visual similarity evaluation, in contrast to image-text aligned ViTs, which are optimized for global semantic alignment. The ablation study also highlights the critical role of our coarse-to-fine strategy for high-resolution images. The performance degrades significantly without both reward components on the ChartMimic and Design2Code benchmarks. Conversely, performance on the lower-resolution UniSVG benchmark is stable, as its images are small enough that our method naturally reduces to a single global comparison, making the coarse and fine-grained rewards equivalent.
@@ 5 Conclusion
In this work, we introduce VinciCoder, a unified multimodal code generation model trained via a two-stage SFT-ViRL strategy. For the SFT stage, we construct a large-scale dataset comprising 1.6M samples from various domains for training. For the subsequent RL stage, we propose a novel ViRL strategy with a coarse-to-fine reward mechanism, which calculates visual similarity rewards from both coarse- and fine-grained image patches. After training on 42k RL data, experimental results across benchmarks demonstrate VinciCoder surpasses contemporary open-source models. More ablation studies validate our proposed ViRL training paradigm and coarse-to-fine reward mechanism.
[…]
@@ 5.2 ChemDraw Metrics
6 Further Experiment Results
@@ 6.1 Comparison with task-specific models
We compare VinciCoder with other domain-specific models, like chart-to-code and image-to-SVG models. The results in Table 6 show that VinciCoder performs comparably to specialized models such as MSRL [chen2025breaking] and ChartMaster [tan2025chartmaster] on the ChartMimic dataset. More domain-specific models are compared in the Appendix. Moreover, its performance on UniSVG exceeds that of fine-tuned specialist models, which validates the effectiveness of our unified framework for multimodal code generation.
@@
ChartCoder [zhao2025chartcoder] 91.4 72.5 74.0
Chart2Code[zhang2025boosting] 62.1 42.9 33.3
MSRL [chen2025breaking] 96.5 78.6 83.8
ChartMaster [tan2025chartmaster] 93.8 78.2 85.1
@@
Table 7: Comparing with task-specific image-to-SVG models. The metrics of finetuned models are from UniSVG [li2025unisvg].
@@ 6.2 ViRL Results
7 Further Analysis
When we generate the refinement data of the image-to-SVG task, we find that directly training a dedicated model does not work well. The training loss is oscillating at 0.5 after training on 360k data for one epoch. We have attempted to train the two more epochs on the previously obtained checkpoint. However, the experimental results indicate that the loss function exhibits persistent oscillations and fails to achieve convergence. To improve the model performance, we propose the visual-based refinement task and directly utilize the checkpoints to generate the data. We visualize the training loss in Figure 9.
Figure 9: Training losses of the first and the next epochs.
8 Refinement Task
We provide the data format of the refinement task in Figure 10. The incomplete code implementation is generated by our trained, dedicated VLM, and the concrete code implementation is from our curated datasets.
@@ 9 Evaluation Prompts
Data Format for Refinement Task Image:
User Instruction: You aim to generate a redrawn code of the given chart. Previously, you generated an incomplete code implementation. Let’s refine this plotting code. It seems to have inaccuracies and isn’t working as expected. Previous generated code: Incomplete Code Implementation Please modify the previous code snippets to ensure the code can render the given image. Assisstant Instruction:
Concrete Code Implementation
Figure 10: Data showcase of the visual-based refinement task.
ChemDraw Benchmark Prompts Instruction: Create a Python script that displays a visual representation of the molecule shown in the image. Utilize ’smiles = <<SMILES>> string’ to define the smile and RKiT or Indigo for plot. Image: Refer to caption